When vector B is added to vector A the resultant is in the negative y direction and has a magnitude of 50N. Given that the x component of vectorA is 15N and the magnitude of vector B is 30N above the negative x axis,find the y component of vector A
1
Expert's answer
2019-01-21T11:17:09-0500

A+B=(0,-50)N.

A_x=15 N,B=30 N

B_x=-A_x=-15 N.

B_y=√(〖30〗^2-(-15)^2 )=15√3 N.

The y component of vector A:

A_y=-50-15√3≈-76 N.